OpenStudy (anonymous):

Negative exponent: 2^-4 + 2^-6/(2^-3) = 5/8?

sam (.sam.):

\[\frac{1}{2^4}+\frac{1}{\frac{2^6}{2^3}}=\frac{5}{8}?\]

OpenStudy (anonymous):

I have the method as to how to answer this question but I'm confused: \[2^{-4}+2^{-6}\div2^{-3} = 1/2^{4}+1/2^{6} \] \[= 2^{2}+1/2^{6}/1/2^{3}\] \[= 2^{2}+1/2^{6}\times(2^{3}/1)\]

OpenStudy (anonymous):

why it suddenly went from 1/2^4 + 1/2^6 to 2^2 + 1/2^6... o.o

sam (.sam.):

Its the same

OpenStudy (anonymous):

but i don't get it.. how it's like that..

sam (.sam.):

2^-4 + 2^-6/(2^-3) = 5/8 2^-4 + 2^-3 = 5/8

sam (.sam.):

You sure your question is correct?

OpenStudy (anonymous):

yes and ohh i get it, thank you!